If I asked you to make 20 ml of .2 M NaOH. How many grams of NaOH would you need to measure out?



1
Expert's answer
2021-04-12T00:04:25-0400

Molar Mass of NaOH = 39.997g/mol

= 39.997 × 2= 79.994

= 79.994 × 20

=1599.88g of NaOH